Calculate the wavelength of a particle with masses 9.11x10-31Kg travelling at speed 1.0x107 m/s and that of a ball of mass 0.1Kg travelling at 35 m/s. Comment on the significance of the wavelengths calculated.

1
Expert's answer
2022-04-20T04:59:04-0400

"\lambda" = h/mv

In this equation, h is Planck’s constant, m is the mass of the particle in kg, and v is the velocity of the particle in m/s. 

"\lambda" = (6.6262×10-34 J s) / (9.11×10-31 kg)(1.0×107 m/s) = 7.28×10-11 m

"\lambda" = (6.6262×10-34 J s) / (0.1 kg)(35 m/s) = 1.89×10-34 m